An independent identically distributed sample size of 3 is chosen from a continuous distribution having density function f(x)=2x,0<x<1, 0 e.w . Find the probability that every member of the sample is less than the median of the distribution.


1
Expert's answer
2021-04-29T17:14:19-0400

The probability of a random sample to be less than median "=\frac{1}{2}" (as 50% values fall below median) therefore probability that all 3 values are below median "=(\frac{1}{2})^3 =\frac{1}{8}"